How many liters each of a 15% acid solution and a 55% acid solution must be used to produce 80 liters of a 50% acid solution? Round to two decimal places if necessary. Help!
Answer by TimothyLamb(4379)

x = liters of 15%
y = liters of 55%
---
x + y = 80
15x + 55y = 50*80
---
put the system of linear equations into standard form
---
x + y = 80
15x + 55y = 4000
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = liters of 15% = 10
y = liters of 55% = 70
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---